LIU-PM-MIN-0323 EDMS 1801164 LIU project team meeting Notes from the meeting held on 4 May 2017 Present: J. Coupard, A. Funken, B. Goddard, K. Hanke, A. Lombardi, M. Meddahi, B. Mikulec, G. Rumolo, R. Scrivens Arising matters (M. Meddahi) No news, we can discuss directly the present LIU budget status with Sylvie’s presentation. LIU spending and EVM status (S. Prodon) Slides are available here. The status was extracted on 1 May. As some EVM progress reports might have been done since then, in a few days the status will be updated as needed. The whole LIU project exhibits a negative cost variance of 400 kCHF and negative schedule variance of 2.6 months (9 MCHF). What is worrying in addition to this large delay appearing only 2 months after the February rebaseling, is the fact that the AC and EV curves have flattened out in April (end of EYETS), while they appeared to proceed with the right slope until March. As of this time of year, the values for the costs should be 19% charged and 57% committed. The fact that for the whole project this is about right is mainly due to the 2.6 MCHF committed on the general management budget. In this sense, we can say that the joker has been already played, but for the rest of the year the WUs will need to commit more and spend more. A first breakdown by the LIU sub-project reveals that the largest schedule variance comes from the PSB (although this is the only project that in 2017 is spending and committing as expected). A breakdown by group is also shown. Some considerations: o BE/BI have planned value equal to actual cost, but low earned value – this causes large cost and schedule variances. It needs to be clarified whether this a reporting problem or the spending is made on something different than what expected for the project – Action: Sylvie to meet Lars/Rhodri o BE/RF have a large schedule variance, plus they need to commit and spend 7 additional MCHF until the end of the year. A meeting has to be organized between Eric, Brennan and Sylvie. Action: Brennan o SMB has a big schedule variance due to PSB, because the item on the B245 is at the moment undefined (the building is finished with money leftover, but since the work still to be done on the building following the POPS accident is not yet defined, the money is kept and translates into a schedule variance). Action: Klaus to clarify the new spening profile and conclude on the specification of the building. o For TE/ABT the schedule variance is real and there is a small commitment budget. Brennan is investigating with his SLs and will come back to the project.. o TE/MSC has recently doubled the schedule variance mainly due to the PSB. LIU project will meet TE/MSC next week. 1 o TE/VSC has a problem with the definition of the jobs with MME. Julie to followup on the document specifying the responsibilities for the Bill Of Materials drawings. In general, the excess money (from items that turned out to be cheaper than predicted) should be immediately moved to the management budgets towards 2019-2020, otherwise it remains in the WUs and appears as very large schedule variances, LIU-ions for PS injectors has a general problem of low commitment, which obviously anticipates a problem of future low spending. Action: Richard to come back in two weeks time with results from internal analysis of the project status. Simulations show that the 2017 target spent money should be reached, but the commitment is lagging dangerously behind (unrealistically below spent money). LIU-PS exhibits large variances, e.g. from BE/BI probably due to EVM reporting and from TE/ABT. Some groups appear to have not only low spending in 2017, but also very low commitment (e.g. RF, ABT, EPC), which means that the money will not be spent in 2017. Simulations show that the 2017 target spent money should be reached, but the slope of the commitment is also too low. LIU-PS: Several groups are contributing to the large schedule variance, i.e. BI, RF, CV, ABT, EPC, VSC. However, the 2017 spending seems quite healthy. The simulations show that by the end of the year the spending could be larger than anticipated, but also here the slope of the commitment needs to pick up again. LIU-SPS: Several groups are contributing to the global schedule variance, in particular BI, RF, STI, ABT, VSC. The main problem for the 2017 spending is with the RF group, which has 2 MCHF committed but 7 MCHF to spend until the end of 2017. Here the simulations indicate that LIU-SPS will spend only half of the anticipated money in 2017. Round table on progress on the various activities Linac4 – A. Lombardi Most of the ongoing work is about planning. The tunnel will be closed by end May. Alessandra is working on making a more detailed planning with precise dates for the RR. End-of-year shutdown needs to be defined yet with Julie. Malika and Alessandra will agree on one of the next LIU-PT meetings in which Alessandra will present the planning. The HST was successfully dismantled and the final celebration took place (Bettina). PSB + PS – K. Hanke PSB: o Last meeting was about the BLM installation status, while the budget follow up is foreseen for the next meeting. o Concerning the document on interlock system, a discussion is ongoing on the situation of beam impact on valves and needs to be finalised. PS: o The question on 80 MHz tuner was discussed at last meeting and it appeared clearly that there is no time now to hire somebody on the simulations, as they 2 need to be carried out in time to allow for the installation of the tuner during one of the 2017 TSs. o On the 11 May there will be a meeting with EN/STI on internal dump. The aim is to make sure that the design is being made with the right specs and to then organize a second review in September to finalise. A discussion on PSB scraper will be kept separated IPAC paper on LIU protons under BE approval. SPS – B. Goddard Conclusions from the aC coating and impedance reduction, many small Nonconformities and problems were found and solved. The technical shielding solution is being reviewed at the end of this month, to allow the launch of the production for the YETS and LS2. Clarified with RF and CO about the uTCA platform, which has been selected to replace VME. CO will support the new platform, prototypes have been ordered for tests of the new LL on the cavities in the lab. RF 200 MHz power upgrade "on track" (although underspending in 2017 needs to be clarified and will be followed up, see action above). An important result is that the SPS started up and is now operating with the new driver amplifiers on the old Philips power plant, which was a crucial step. Progressing well on all other technical items. Ions PS Injectors – R. Scrivens No news. Planning and Installation – J. Coupard Linac4 work: one issue is the debuncher installation (it cannot be connected with the waveguide for lack of space in tunnel) – this will be reviewed next Tuesday to check what work is needed for a definitive solution. This will not impact the RR, as the debuncher will only be powered later. Spending linked to this item will be on operational budget. Realignment of the 3 MeV part will be assessed this week. A point on Linac4 connection will be made at the next LIU-PSB meeting. Three ECRs will be sent out by end of the month. There was an integration meeting yesterday and the notes have been sent around. S. Deleval presented the CV work of PS, for which the ECR will be sent as soon as the integration is validated. The organization of the integration meetings for LIU-SPS is still being finalized. List of non-approved PLAN activities still has to be sent to the group leaders (as was decided at the last LIU-PT meeting). Safety – A. Funken New mechanical safety rule GSI-M6 for machinery: the evaluation of the new wire scanners was made with HSE and BE-BI last week. Mostly the risk analysis needs to be finalized and existing procedure for intervention to be attached. LIU-SPS Ring and TF lines descriptive part: Anne will integrate the last contributions Brennan received for launching the EDMS review (by contributors) 3 tomorrow or beginning of next week. Still missing (minor) information that will be collected during the EDMS process. Official homologation of Linac4 safety file (general safety an RP parts) received on 26th of April. To be noted that requests are to be considered for other safety files to be presented to ASN/OFSP. Action: Anne to send the document to the LIU-PT. Alessandra added that within one year, the estimate of the radioactive waste produced during Linac4 operation has to be prepared and sent to ASN/OFSP. Action: Alessandra/Anne to discuss further this topic for defining how to proceed. AOB The next LIU-PT meeting is scheduled on Thursday 11th May 2017. Minutes by G. Rumolo 4
© Copyright 2025 Paperzz